# Lipschitz Bandits with Stochastic Delayed Feedback

### v1

@@ 5 Delay with Unbounded Support: Delayed Lipschitz Phased Pruning
In practice, the bounded delay assumption may not always hold, as feedback can be missing or censored, corresponding to an infinite delay. To address this challenge, and inspired by the success of elimination-based methods in Lipschitz bandits (feng2022lipschitz, ; kang2023robust, ), we propose a novel phased learning algorithm called Delayed Lipschitz Phased Pruning (DLPP). DLPP accumulates reliable feedback over carefully scheduled intervals and achieves a regret bound that is nearly optimal up to logarithmic factors, as supported by our lower bound analysis. While BLiN (feng2022lipschitz, ) is the first elimination-based algorithm for Lipschitz bandits, our approach is fundamentally different in both methodology and analysis. The differences in DLPP from BLiN are summarized as below:
@@
10:    If $v(C)\geq v_{m}=(4\log T+2\log(2/\delta))/r_{m}^{2}$, remove $C$ from $\mathcal{B}_{m}^{+}$.
@@
In each phase $m$, our algorithm proceeds as follows. Let $\mathcal{B}_{m}$ be the set of currect active balls, and set the active set $\mathcal{B}_{m}^{+}=\mathcal{B}_{m}$. For each ball $B\in\mathcal{B}_{m}^{+}$, we sample an arm $x\in B$ uniformly and play it, schedule the reward to be observed at time $t+\tau_{t}$. Then we observe any incoming scheduled reward $(C,y_{C})$ and update the corresponding empirical average reward $\hat{\mu}(C)$ and number of observed times $v(C)$ of $C$. If $v(C)\geq v_{m}=(4\log T+2\log(2/\delta))/r_{m}^{2}$, we remove $C$ from the active set $\mathcal{B}_{m}^{+}$, and do not play it anymore in this phase (i.e. line 6 of Algorithm 2). Increment the timer $t=t+1$ and exit the algorithm if $t>T$. The number $v_{m}$ is carefully designed such that the empirical average reward over a ball concentrated over its true mean with high probability. We assume that tuple $(C,y_{C})$ is observed as feedback. It is necessary to know which ball an arm $x$ belongs to, as the balls may overlap; however, identifying the specific arm within a ball is not required, since all arms in a ball are treated uniformly to represent the ball and sampled at random. The procedure proceeds in a Round-Robin fashion such that we eliminate the balls with low empirical average reward in the next step.
@@
Consider an instance $(\mathcal{A},\mathcal{D},\mu)$ of the delayed Lipschitz Bandit problems with time horizon $T$ and a delay distribution with quantile function $Q(p)$, with probability at least $1-\delta$, the Algorithm 2 attains regret
@@
The claimed bound is achieved minimizing over a single quantile $p\in(0,1]$, in the sense that it reaches the smallest order. When there is no delay, i.e. $\tau_{max}=0$, the regret bound reduces to $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}\right)$, which is exactly the regret bound of Lipschitz bandits; this is the same as in Theorem 1. The increase in regret due to delays in the claimed bound does not scale with the underlying zooming dimension $d_{z}$. Take $p=0.5$, since $\tau_{med}=Q(0.5)$ denotes the median of the delays, the bound becomes $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{med}\right)$, which is essentially the same with the order of Theorem 1. This quantile-based bound provides a flexible characterization of the delay distribution’s effect on regret, allowing the algorithm to adapt to the effective central mas of delays rather than worst-case outliers. To see why the increase in regret due to delays should scale with a certain quantile, one can simulate a black box algorithm for the rounds that delay is smaller than $\tau_{med}$, and take the last action of the black box algorithm for the rest rounds (approximately half of the rounds). Since the rewards are stochastic and independent of both time and delay, the regret incurred on rounds with delays greater than $\tau_{med}$ is comparable to the regret of the black-box algorithm on the remaining rounds, as a result, the total regret is essentially twice that of the black-box algorithm (lancewicki2021stochastic, ).
@@ 6 Lower Bound
The lower bound is proved using a specified delay distribution such that the delay is a fixed value $\tau_{0}$ with probability $p$, and $\infty$ otherwise (missing feedback). The first term in Eq. (3) retains the optimal regret rate of Lipschitz bandits, up to logarithmic factors, scaling with the delay quantiles. The first term also matches the existing lower bound, proved through a reduction from the non-delayed lower bound for Lipschitz bandits (kleinberg2019bandits, ; slivkins2011contextual, ). The core idea is to simulate a delayed variant of a Lipschitz bandit algorithm by introducing a Bernoulli sampling mechanism with parameter $p$. At each round, a Bernoulli random variable is sampled; if its outcome is 1, the original Lipschitz bandit algorithm selects the action, thereby mirroring the behavior of its delayed counterpart. Otherwise, only the delayed version takes an action. This probabilistic coupling ensures that the original and delayed algorithms align with probability $p$, enabling a controlled simulation of delay effects. Compared to the upper bound in Theorem 3, there is an additional term. Specifically, the last term in Eq. (3) arises from the fact that the algorithm receives no feedback during the first $\tau_{0}=Q(p)$ rounds. As a result, the learner is unable to distinguish between arms in this initial period, and there exists at least one problem instance in which the learner incurs a regret of $\tau_{0}\bar{\Delta}$ where $\bar{\Delta}=\int_{\mathcal{A}}\Delta(x)/\int_{\mathcal{A}}1$ per round over the first $\tau_{0}$ rounds. This demonstrates that the lower bound essentially matches the upper bound, implying that our regret bound is nearly optimal up to logarithmic factors. A more detailed discussion and complete proof are provided in Appendix C.
@@ 7 Experiments
In this section, we present simulation results to demonstrate the performance of our proposed algorithms under various delay conditions. To the best of our knowledge, there are no existing algorithms that directly address the Lipschitz bandit problem with stochastic delayed feedback. Although the BLiN algorithm proposed in feng2022lipschitz shares some structural similarities with our DLPP algorithm, it is not applicable in our setting. In particular, BLiN relies on immediate or batched feedback, whereas in our model, the presence of additional delays and the absence of a batched communication constraint render region-level reward estimates unreliable.
To thoroughly validate the robustness of our proposed methods, we employ three types of expected reward functions under both bounded and unbounded delay conditions, each with varying average delays. We first consider the metric space $([0,1],|\cdot|)$ with two expected reward functions that behave differently around their maxima, including a triangle function $\mu(x)=0.8-0.9\cdot|x-0.4|$, and a sine wave function $\mu(x)=\frac{2}{3}\left|\sin\left(\frac{5\pi}{3}x\right)\right|$ with two different maximum point ($x=0.3,0.9$). We further consider the more complicated arm metric space $([0,1]^{2},\|\cdot\|_{\infty})$, and the expected reward function is a two dimensional function $\mu(x)=1-0.7\left\|{x-x_{1}}\right\|_{2}-0.4\left\|{x-x_{2}}\right\|_{2},$ where $x_{1}=(0.7,0.8)$ and $x_{2}=(0,0.1)$, with $\mu^{*}=\mu(x_{1})$. We set with time horizon $T=60,000$ and false probability rate $\delta=0.01$, and we repeat the experiment over independent trials $B=30$ and take the average cumulative regret. We evaluate the algorithms with different delay distributions. In the first case, we consider the uniform distribution for the bounded delay distributions, whereas in the second case, we implement the geometric distribution for the unbounded delays. As in the empirical study done by chapelle2014modeling , delays are empirically shown to have an exponential decay. We consider both cases with different mean that $\mathbb{E}[\tau]\in\{20,50\}$. Since there is no existing delayed Lipschitz bandit algorithm, we also include the result of the delay-free setting as the baseline to compare the performance. The graphic regret curves are reported in Figure 1.
@@ 8 Conclusion
In this work, we introduce the problem of Lipschitz bandits under stochastic delayed feedback and propose two algorithms that address bounded and unbounded delays, respectively, supported by comprehensive theoretical analysis. For the bounded-delay setting, we propose a delay-aware zooming algorithm called Delayed Zooming Algorithm that matches the optimal performance of the delay-free case (kleinberg2008multi, ) up to an additive term that scales with the maximal delay. For unbounded delays, we propose Delayed Lipschitz Phased Pruning (DLPP) that novelly accumulates reliable feedback over carefully scheduled intervals and achieves near-optimal regret bounds, with an additional term that depends on quantiles of the delay distribution. We further establish a lower bound that nearly matches our deduced upper bounds up to logarithmic factors, demonstrating that the regret bound of DLPP is nearly optimal. The effectiveness of our proposed algorithms is finally validated through numerical experiments.
@@ A.2 Proof of Theorem 1
We follow the proof idea of [12] and [5]. Fix round $t$, let $V_{t}$ be the set of all arms that are active at time $t$, and let
@@ B.1 Some Useful Lemmas
$\mathcal{E}_{1}:=\left\{|\hat{\mu}_{m}(B)-\mu(x)|\leq r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\forall x\in B,\forall B\in\mathcal{B}_{m},\forall 1\leq m\leq m^{*}\right\},$
@@
For a fixed ball $B\in\mathcal{B}_{m}$ that is removed from $\mathcal{B}_{m}^{+}$, there are $v_{m}=(4\log T+2\log(2/\delta))/r_{m}^{2}$ observations from $B$ has been observed, thus the empirical average reward of $B$ and its expectation is
@@
$|\mathbb{E}[\hat{\mu}_{m}(B)]-\mu(x)|=\left|\frac{1}{v_{m}}\sum_{i=1}^{v_{m}}(\mu(x_{B,i})-\mu(x))\right|\leq r_{m}.$
@@
$\Pr\left(|\hat{\mu}_{m}(B)-\mu(x)|\leq r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\quad\forall x\in B\right)\geq 1-\frac{\delta}{T^{2}}.$
@@
$\Pr\left(|\hat{\mu}_{m}(B)-\mu(x)|\leq r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\quad\forall x\in B,\forall B\in\mathcal{B}_{m}\right)\geq 1-\frac{\delta}{T}.$
@@
$\hat{\mu}_{m}(B)-\mu(x)\leq r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},$
@@
$\mu(x^{*})-\hat{\mu}_{m}(B^{*})\leq r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}}.$
@@
For $m=1$, it hold trivially that $8r_{m-1}=8>1\geq\Delta(x)$. For $2\leq m\leq m^{*}+1$, we consider the previous complete phase m1m-1 so that the lemma also holds for the incomplete phase $m^{*}+1$. Suppose for any ball $B\in\mathcal{B}_{m}$ and $x\in B,x^{*}\in B_{m}^{*}$, in phase m1m-1, $x\in B_{0}\in\mathcal{B}_{m-1}$, where $B_{0}$ is the parent ball of $B$ such that it contains $x$ in phase m1m-1, and let $x^{*}\in B_{m-1}^{*}\in\mathcal{B}_{m-1}$ since the optimal arm is not pruned. By Lemma 11, for any $B\in\mathcal{B}_{m}$ and any $x\in B$
$\Delta(x)$ $=\mu^{*}-\mu(x)\leq\hat{\mu}_{m-1}(B_{m-1}^{*})-\hat{\mu}_{m-1}(B_{0})+2r_{m-1}+2\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m-1}}}.$
@@ B.2 Proof of Theorem 3
We modify the proof in [14] by relating $n_{m}$ and $v_{m}$ using previous lemmas to show that the addtional term in the regret bound incurred by the delays scales with quantiles.
@@
Fix $p\in(0,1]$, and define $\mathcal{E}=\mathcal{E}_{1}\cap\mathcal{E}_{2}$ as in Lemma 1213, it holds that $\Pr[\mathcal{E}]\geq 1-2\delta$. For any phase $m$ and any ball $B\in\mathcal{B}_{m}$, let $s_{B}$ be the last time that ball $B$ is played, thus we have $v_{s_{B}-1}(B)<v_{m}=(4\log T+2\log(2/\delta))/r_{m}^{2}$. Under the clean event $\mathcal{E}_{2}$, at least one of the two following event is true:
$n_{s_{B}-Q(p)-1}(B)\leq\frac{2}{p}v_{s_{B}-1}(B)\leq\dfrac{8\log T+4\log(2/\delta)}{pr_{m}^{2}},$

### v2

@@ 5 Delay with Unbounded Support: Delayed Lipschitz Phased Pruning
In practice, the bounded delay assumption may not always hold, as feedback can be missing or censored, corresponding to an infinite delay. To address this challenge, and inspired by the success of elimination-based methods in Lipschitz bandits (Feng et al., 2022; Kang et al., 2023), we propose a novel phased learning algorithm called Delayed Lipschitz Phased Pruning (DLPP). DLPP accumulates reliable feedback over carefully scheduled intervals and achieves a regret bound that is nearly optimal up to logarithmic factors, as supported by our lower bound analysis. While BLiN (Feng et al., 2022) is the first elimination-based algorithm for Lipschitz bandits, our approach is fundamentally different in both methodology and analysis. The differences in DLPP from BLiN are summarized as below:
@@
10:    If $v(C)\geq v_{m}=(2\log T+\log(2/\delta))/2r_{m}^{2}$, remove $C$ from $\mathcal{B}_{m}^{+}$.
@@
In each phase $m$, our algorithm proceeds as follows. Let $\mathcal{B}_{m}$ be the set of currect active balls, and set the active set $\mathcal{B}_{m}^{+}=\mathcal{B}_{m}$. For each ball $B\in\mathcal{B}_{m}^{+}$, we sample an arm $x\in B$ uniformly and play it, schedule the reward to be observed at time $t+\tau_{t}$. Then we observe any incoming scheduled reward $(C,y_{C})$ and update the corresponding empirical average reward $\hat{\mu}(C)$ and number of observed times $v(C)$ of $C$. If $v(C)\geq v_{m}=(2\log T+\log(2/\delta))/2r_{m}^{2}$, we remove $C$ from the active set $\mathcal{B}_{m}^{+}$, and do not play it anymore in this phase (i.e. line 6 of Algorithm 2). Increment the timer $t=t+1$ and exit the algorithm if $t>T$. The number $v_{m}$ is carefully designed such that the empirical average reward over a ball concentrated over its true mean with high probability. We assume that tuple $(C,y_{C})$ is observed as feedback. It is necessary to know which ball an arm $x$ belongs to, as the balls may overlap; however, identifying the specific arm within a ball is not required, since all arms in a ball are treated uniformly to represent the ball and sampled at random. The procedure proceeds in a Round-Robin fashion such that we eliminate the balls with low empirical average reward in the next step.
@@
Consider an instance $(\mathcal{A},\mathcal{D},\mu)$ of the delayed Lipschitz Bandit problems with time horizon $T$ and a delay distribution with quantile function $Q(p)$, with probability at least $1-2\delta$, the Algorithm 2 attains regret
@@
The claimed bound is achieved minimizing over a single quantile $p\in(0,1]$, in the sense that it reaches the smallest order. When there is no delay, i.e. $\tau_{max}=0$, the regret bound reduces to $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}\right)$, which is exactly the regret bound of Lipschitz bandits; this is the same as in Theorem 1. The increase in regret due to delays in the claimed bound does not scale with the underlying zooming dimension $d_{z}$. Take $p=0.5$, since $\tau_{med}=Q(0.5)$ denotes the median of the delays, the bound becomes $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{med}\right)$, which is essentially the same with the order of Theorem 1. This quantile-based bound provides a flexible characterization of the delay distribution’s effect on regret, allowing the algorithm to adapt to the effective central mas of delays rather than worst-case outliers. To see why the increase in regret due to delays should scale with a certain quantile, one can simulate a black box algorithm for the rounds that delay is smaller than $\tau_{med}$, and take the last action of the black box algorithm for the rest rounds (approximately half of the rounds). Since the rewards are stochastic and independent of both time and delay, the regret incurred on rounds with delays greater than $\tau_{med}$ is comparable to the regret of the black-box algorithm on the remaining rounds, as a result, the total regret is essentially twice that of the black-box algorithm (Lancewicki et al., 2021).
@@ 6 Lower Bound
The lower bound is proved using a specified delay distribution such that the delay is a fixed value $\tau_{0}$ with probability $p$, and $\infty$ otherwise (missing feedback). The first term in Eq. (3) retains the optimal regret rate of Lipschitz bandits, up to logarithmic factors, scaling with the delay quantiles. The first term also matches the existing lower bound, proved through a reduction from the non-delayed lower bound for Lipschitz bandits (Kleinberg et al., 2019; Slivkins, 2011). The core idea is to simulate a delayed variant of a Lipschitz bandit algorithm by introducing a Bernoulli sampling mechanism with parameter $p$. At each round, a Bernoulli random variable is sampled; if its outcome is 1, the original Lipschitz bandit algorithm selects the action, thereby mirroring the behavior of its delayed counterpart. Otherwise, only the delayed version takes an action. This probabilistic coupling ensures that the original and delayed algorithms align with probability $p$, enabling a controlled simulation of delay effects. Compared to the upper bound in Theorem 3, there is an additional term. Specifically, the last term in Eq. (3) arises from the fact that the algorithm receives no feedback during the first $\tau_{0}=Q(p)$ rounds. As a result, the learner is unable to distinguish between arms in this initial period, and there exists at least one problem instance in which the learner incurs a regret of $\tau_{0}\bar{\Delta}$ where $\bar{\Delta}=\int_{\mathcal{A}}\Delta(x)/\int_{\mathcal{A}}1$ per round over the first $\tau_{0}$ rounds. This demonstrates that the lower bound essentially matches the upper bound, implying that our regret bound is nearly optimal up to logarithmic factors. A more detailed discussion and complete proof are provided in Appendix C.
@@ 7 Experiments
In this section, we present simulation results to demonstrate the performance of our proposed algorithms under various delay conditions. To the best of our knowledge, there are no existing algorithms that directly address the Lipschitz bandit problem with stochastic delayed feedback. Although the BLiN algorithm proposed in Feng et al. (2022) shares some structural similarities with our DLPP algorithm, it is not applicable in our setting. In particular, BLiN relies on immediate or batched feedback, whereas in our model, the presence of additional delays and the absence of a batched communication constraint render region-level reward estimates unreliable.
To thoroughly validate the robustness of our proposed methods, we employ three types of expected reward functions under both bounded and unbounded delay conditions, each with varying average delays. We first consider the metric space $([0,1],|\cdot|)$ with two expected reward functions that behave differently around their maxima, including a triangle function $\mu(x)=0.8-0.9\cdot|x-0.4|$, and a sine wave function $\mu(x)=\frac{2}{3}\left|\sin\left(\frac{5\pi}{3}x\right)\right|$ with two different maximum point ($x=0.3,0.9$). We further consider the more complicated arm metric space $([0,1]^{2},\|\cdot\|_{\infty})$, and the expected reward function is a two dimensional function $\mu(x)=1-0.7\left\|{x-x_{1}}\right\|_{2}-0.4\left\|{x-x_{2}}\right\|_{2},$ where $x_{1}=(0.7,0.8)$ and $x_{2}=(0,0.1)$, with $\mu^{*}=\mu(x_{1})$. We set with time horizon $T=60,000$ and false probability rate $\delta=0.01$, and we repeat the experiment over independent trials $B=30$ and take the average cumulative regret. We evaluate the algorithms with different delay distributions. In the first case, we consider the uniform distribution for the bounded delay distributions, whereas in the second case, we implement the geometric distribution for the unbounded delays. As in the empirical study done by Chapelle (2014), delays are empirically shown to have an exponential decay. We consider both cases with different mean that $\mathbb{E}[\tau]\in\{20,50\}$. Since there is no existing delayed Lipschitz bandit algorithm, we also include the result of the delay-free setting as the baseline to compare the performance. The graphic regret curves are reported in Figure 1.
@@ 8 Conclusion
In this work, we introduce the problem of Lipschitz bandits under stochastic delayed feedback and propose two algorithms that address bounded and unbounded delays, respectively, supported by comprehensive theoretical analysis. For the bounded-delay setting, we propose a delay-aware zooming algorithm called Delayed Zooming Algorithm that matches the optimal performance of the delay-free case (Kleinberg et al., 2008) up to an additive term that scales with the maximal delay. For unbounded delays, we propose Delayed Lipschitz Phased Pruning (DLPP) that novelly accumulates reliable feedback over carefully scheduled intervals and achieves near-optimal regret bounds, with an additional term that depends on quantiles of the delay distribution. We further establish a lower bound that nearly matches our deduced upper bounds up to logarithmic factors, demonstrating that the regret bound of DLPP is nearly optimal. The effectiveness of our proposed algorithms is finally validated through numerical experiments.
@@
Acknowledgements
We truly appreciate the constructive feedback from the anonymous reviewers and area chair. This work was partially supported by the National Science Foundation under grants DMS-2210388 and DMS-2515304 GFI.
@@ A.2 Proof of Theorem 1
We follow the proof idea of Kleinberg et al. (2019) and Joulani et al. (2013). Fix round $t$, let $V_{t}$ be the set of all arms that are active at time $t$, and let
@@ B.1 Some Useful Lemmas
$\mathcal{E}_{1}:=\left\{|\hat{\mu}_{m}(B)-\mu(x)|\leq 2r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\forall x\in B,\forall B\in\mathcal{B}_{m},\forall 1\leq m\leq m^{*}\right\},$
@@
For a fixed ball $B\in\mathcal{B}_{m}$ that is removed from $\mathcal{B}_{m}^{+}$, there are $v_{m}=(2\log T+\log(2/\delta))/2r_{m}^{2}$ observations from $B$ has been observed, thus the empirical average reward of $B$ and its expectation is
@@
$|\mathbb{E}[\hat{\mu}_{m}(B)]-\mu(x)|=\left|\frac{1}{v_{m}}\sum_{i=1}^{v_{m}}(\mu(x_{B,i})-\mu(x))\right|\leq 2r_{m}.$
@@
$\Pr\left(|\hat{\mu}_{m}(B)-\mu(x)|\leq 2r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\quad\forall x\in B\right)\geq 1-\frac{\delta}{T^{2}}.$
@@
$\Pr\left(|\hat{\mu}_{m}(B)-\mu(x)|\leq 2r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},\quad\forall x\in B,\forall B\in\mathcal{B}_{m}\right)\geq 1-\frac{\delta}{T}.$
@@
$\hat{\mu}_{m}(B)-\mu(x)\leq 2r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}},$
@@
$\mu(x^{*})-\hat{\mu}_{m}(B^{*})\leq 2r_{m}+\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m}}}.$
@@
For $m=1$, it hold trivially that $16r_{m-1}=16>1\geq\Delta(x)$. For $2\leq m\leq m^{*}+1$, we consider the previous complete phase m1m-1 so that the lemma also holds for the incomplete phase $m^{*}+1$. Suppose for any ball $B\in\mathcal{B}_{m}$ and $x\in B,x^{*}\in B_{m}^{*}$, in phase m1m-1, $x\in B_{0}\in\mathcal{B}_{m-1}$, where $B_{0}$ is the parent ball of $B$ such that it contains $x$ in phase m1m-1, and let $x^{*}\in B_{m-1}^{*}\in\mathcal{B}_{m-1}$ since the optimal arm is not pruned. By Lemma 11, for any $B\in\mathcal{B}_{m}$ and any $x\in B$
$\Delta(x)$ $=\mu^{*}-\mu(x)\leq\hat{\mu}_{m-1}(B_{m-1}^{*})-\hat{\mu}_{m-1}(B_{0})+4r_{m-1}+2\sqrt{\dfrac{4\log T+2\log(2/\delta)}{v_{m-1}}}.$
@@ B.2 Proof of Theorem 3
We modify the proof in Feng et al. (2022) by relating $n_{m}$ and $v_{m}$ using previous lemmas to show that the addtional term in the regret bound incurred by the delays scales with quantiles.
@@
Fix $p\in(0,1]$, and define $\mathcal{E}=\mathcal{E}_{1}\cap\mathcal{E}_{2}$ as in Lemma 1213, it holds that $\Pr[\mathcal{E}]\geq 1-2\delta$. For any phase $m$ and any ball $B\in\mathcal{B}_{m}$, let $s_{B}$ be the last time that ball $B$ is played, thus we have $v_{s_{B}-1}(B)<v_{m}=(2\log T+\log(2/\delta))/2r_{m}^{2}$. Under the clean event $\mathcal{E}_{2}$, at least one of the two following event is true:
$n_{s_{B}-Q(p)-1}(B)\leq\frac{2}{p}v_{s_{B}-1}(B)\leq\dfrac{2\log T+\log(2/\delta)}{pr_{m}^{2}},$